OpenStudy (anonymous):

a canon shell is launched from a cliff 400m above a plain at a velocity of 250ms at 60 above the horizontal. 1. initial ux=216.5ms uy=125 2.maximum height is 411.046 m 3.total time of flight is 31.25sec 4. range is 3906.25m 5. the impact velocity of the shell ? how do i find this?

OpenStudy (emmigrace222):

i have no idea im sorry

OpenStudy (atlas):

i hope you know the basic equations of motion. The trick is to find the horizontal and vertical components of velocity separately and apply the equations of motion on them

OpenStudy (anonymous):

yea i am doing that i am not finding a bigger angle

OpenStudy (anonymous):

than 60'

OpenStudy (atlas):

i think 60 is the height isn't it

OpenStudy (anonymous):

noo its an angle height is 400 m but total height is 411.046

OpenStudy (atlas):

one simple way that I can think of: u know ux will remain unchanged as there is no acceleration in the x direction maximum height is given and at maximum height uy =0

OpenStudy (atlas):

so you can find vy (velocity in y direction at the point of impact) using v^2 = u^2 + 2as

OpenStudy (atlas):

u already know vx (=ux) and you will find vy you can easily find the velocity at the point of impact

OpenStudy (atlas):

you can additionally find the angle at which it strikes the ground if you would like to

OpenStudy (anonymous):

uy=216.5 ay=-9.8ms t =31.25 so vy = uy +at vy= 216.5+ (-9.8) 31.25 = -89.75ms

OpenStudy (atlas):

no

OpenStudy (atlas):

use v^2 = u^2 +2 as where s is the max. height

OpenStudy (atlas):

and uy =0 at max height

OpenStudy (atlas):

vy^2 = 2 (9.8) *411.046

OpenStudy (anonymous):

the same answer is coming

OpenStudy (anonymous):

see i wanna find an angle tan -1 ( 89.75/125) which shuld be bigger than 60

OpenStudy (atlas):

oh yes you will get the same

OpenStudy (atlas):

and both are right

OpenStudy (atlas):

it gives 112 degrees i gues

OpenStudy (anonymous):

no 35

OpenStudy (atlas):

let me check

OpenStudy (atlas):

yeah right 35......but why do you say the angle should be greater than 60

OpenStudy (anonymous):

i dunno my teacher said that! -_-

OpenStudy (anonymous):

i ll check with him tomorrow thnx for the help !

OpenStudy (atlas):

first of all i believe that 60 above the horizontal cannot be mentioning the angle but distance above the horizon

OpenStudy (atlas):

u can check for yourself by finding the angle by tan-1(uy/ux)